A sample of 200 persons is asked about their handedness. A two way table of observed counts follow
1
Expert's answer
2020-04-10T17:21:08-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
 & Left-handed & Right-handed& Total\\ \hline
 Men & 11 &79 & 90 \\
 \hdashline
 Women & 9 & 101 & 110 \\
\hdashline
Total & 20 & 180 & 200
\end{array}"

Let "M=" selected person is a men, "W=" selected person is a women, "L=" selected person is left-handed, "R=" selected person is right-handed.

If one person is randomly selected

a. "P(W)=\dfrac{110}{200}=\dfrac{11}{20}=0.55"


b. "P(R)=\dfrac{180}{200}=\dfrac{9}{10}=0.9"


c. "P(M\cap R)=\dfrac{79}{200}=0.395"


d. "P(W\cup L)=\dfrac{20+110-9}{200}=\dfrac{121}{200}=0.605"


e. "P(M|L)=\dfrac{11}{20}=0.55"


f. "P(R|W)=\dfrac{101}{110}\approx0.9182"


If one person is randomly selected with replacement


"P(first\ M\cap L, \ second\ M\cap R )=\dfrac{11}{200}\cdot\dfrac{79}{200}=""=\dfrac{869}{40000}=0.021725"

"P(first\ W\cap L, \ second\ W\cap L )=\dfrac{9}{200}\cdot\dfrac{9}{200}=""=\dfrac{81}{40000}=0.002025"

If one person is randomly selected without replacement


"P(first\ M\cap L, \ second\ M\cap R )=\dfrac{11}{200}\cdot\dfrac{79}{199}=""=\dfrac{869}{39800}\approx0.021834"


"P(first\ W\cap L, \ second\ W\cap L )=\dfrac{9}{200}\cdot\dfrac{9}{199}=""=\dfrac{81}{39800}\approx0.002035"